# Workflow as a Service Broker in Cloud Environment: A Systematic Literature Review

### v1

@@ 4.2.3. Task Concurrency
Task concurrency defines the number of tasks that can be executed concurrently on the same resource. Most studies adopt a single mode, where only one task is scheduled on each resource at a time (Chen et al., 2021; Taghinezhad-Niar and Taheri, 2023). However, some studies explore the multiple mode, which permits the concurrent execution of several tasks on the same resource. For instance, in (Taghavi et al., 2023; Saeedizade and Ashtiani, 2021), brokers rent large virtual machines (VMs) with a high number of CPU cores and significant memory, bundling tasks together and scheduling them on these VMs as containers. Similarly, Silva et al. (Silva et al., 2024) utilize multi-threading on a single node to enable the simultaneous execution of multiple tasks.
4.2.4. Data Transmission
@@
Data transmission outlines how output data from parent (predecessor) tasks is delivered to child (successor) tasks. In the direct (or peer-to-peer) approach, data is transferred straight from the resource executing the predecessor task to the resource assigned to the successor task. This method requires synchronous communication, meaning the successor task must already be scheduled, and its resource allocation must be determined beforehand. A notable advantage of this approach is that data transfer time between tasks scheduled on the same resource becomes negligible. However, it may not be ideal for scheduling strategies that delay the execution of successor tasks. In contrast, the indirect (or shared) method involves the predecessor task storing its output data in shared storage, such as Amazon S3 or Amazon EFS. The successor task can then retrieve this data asynchronously, enabling more flexibility in task scheduling. Researchers have utilized both direct (Chen et al., 2018; Arabnejad et al., 2019) and indirect (Rajasekar and Palanichamy, 2022, 2021) methods in their studies.
4.2.5. Scheduling Criteria
This sub-section explores the scheduling criteria frequently used as optimization objectives or constraints in the MWSP. The most common criteria are the execution time of workflows and the rental cost of resources, which are addressed in almost all studies, either as optimization goals or constraints (such as deadlines or budgets). In addition to these, some studies focus on maximizing resource utilization (Liu et al., 2024; Chen et al., 2016; Jiang et al., 2011). Another key objective is minimizing energy consumption, which is generally a provider-focused goal. Although it is rarely considered a client- or broker-side concern, some studies suggest that brokers could reduce energy usage by adjusting the CPU frequencies of leased resources (Taghinezhad-Niar and Taheri, 2023; Taghinezhad-Niar et al., 2022). In hybrid cloud environments, where both private and public clouds are used, researchers often aim to reduce energy consumption within the private cloud while disregarding energy usage in the public cloud, as it is not typically a user concern (Li et al., 2022d; Sun et al., 2023).
@@ 4.2.5. Scheduling Criteria
Fairness is another criterion considered in some studies, although its definition varies. Some works, such as (Chen et al., 2021; Wu and Yuandou, 2018), approach fairness as a resource-centric measure to ensure balanced workload distribution across resources. In contrast, others, including (Li et al., 2024b, c; Rezaeian et al., 2019), define fairness as a user-centric measure that focuses on equitable resource allocation among workflows. For example, (Li et al., 2024b) introduces the concept of a workflow’s loss rate, which combines two factors: slowdown (the ratio of a workflow’s completion time when scheduled alongside others to its completion time when scheduled alone) and overspending (the cost ratio under the same conditions). Fairness is then defined as minimizing the disparity in loss rates across all workflows. Additional objectives considered by researchers include improving reliability (Taghinezhad-Niar and Taheri, 2023; Alam et al., 2024; Sampaio and Barbosa, 2020), ensuring privacy (Sun et al., 2023; Sharif et al., 2014), and enhancing security (Alam et al., 2024; Roy et al., 2024).
4.2.6. Scheduling Type
@@
Workflow scheduling can be classified into two primary types: static and dynamic (Rodriguez and Buyya, 2017). Static scheduling involves assigning tasks to resources prior to the start of workflow execution. This approach typically produces high-quality solutions since it considers crucial structural aspects of the workflow, such as the critical path. However, static scheduling is limited by its reliance on precise runtime estimations and its inability to adapt to unforeseen changes in resource availability, making it less suitable for the uncertain nature of cloud environments. On the other hand, dynamic scheduling makes decisions during runtime, assigning tasks as they become ready—that is, once all their predecessor tasks have been completed. This enables dynamic methods to respond effectively to changes and uncertainties in the cloud environment. Nevertheless, these algorithms often make greedy, task-specific decisions, lacking a global perspective on the entire workflow. Consequently, the solutions generated by dynamic approaches are generally less optimal compared to those of static methods. To address the shortcomings of both approaches, hybrid strategies have been proposed. These methods begin by creating a static execution plan for the workflow and then apply dynamic adjustments during runtime to account for environmental uncertainties. This combination leverages the strengths of both static and dynamic scheduling to produce more robust solutions.
@@
In static multiple workflow scheduling, two primary approaches have been used: one-by-one and combined. In the one-by-one method, workflows are scheduled sequentially, based on priority, while considering resource availability. For example, Rezaeian et al. (Rezaeian et al., 2019) apply the one-by-one approach using four different workflow selection methods: Ordered, which schedules workflows based on a predefined priority; round-robin, which iterates over workflows, scheduling one task from each workflow at a time; weighted round-robin, where workflows are selected according to a weighted formula before scheduling a task; and unfairness heuristic, which selects workflows based on their level of unfairness. Rimal and Maier (Rimal and Maier, 2017) employ three different workflow selection strategies: FCFS (First Come First Serve), EASY Backfilling, and MCT (Minimum Completion Time). In contrast, the combined approach merges all workflows into a single large workflow before scheduling it on available resources. References (Alam et al., 2024; Ahmad et al., 2022) use a similar approach, where they break down each workflow into levels based on task depth. Then, they create partitions where each partition contains tasks at the same level across different workflows. Other studies merge workflows by adding a dummy entry task to connect the input workflows’ actual entry tasks, and a dummy exit task to connect the workflows’ exit tasks (Taheri et al., 2023; Qin and Shao, 2025; Zheng et al., 2019).
@@
Two strategies are commonly used for dynamic multiple workflow scheduling: immediate and periodic. In the immediate strategy, tasks are scheduled as soon as they become ready. However, this approach faces two challenges: first, the runtime overhead caused by executing the scheduling algorithm for each task as it becomes ready, and second, the greedy decision-making process for each task without considering the future tasks and their potential impact on the current scheduling. To address these issues, the periodic strategy was introduced. This method divides time into intervals, collecting all ready tasks within each interval and scheduling them all at once at the end of the interval. While this approach resolves the problems associated with immediate scheduling, it also introduces a delay, causing ready tasks to wait until the end of the interval, potentially missing deadlines. Most previous work in dynamic scheduling adopts the immediate strategy (Liu et al., 2021; Ma et al., 2021; Rajasekar and Palanichamy, 2022), although some utilize the periodic strategy (Taghavi et al., 2023; Lin et al., 2024). Chen et al. (Chen et al., 2017) introduced a hybrid strategy, where they use periodic scheduling but select tasks at the start of each interval that are predicted to have their earliest start time within the current interval. These tasks are scheduled on available resources ahead of time.
@@
Only two studies have employed a combination of both static and dynamic scheduling types. Taheri et al. (Taheri et al., 2023) proposed a framework consisting of two phases: static planning and dynamic scheduling. In the static planning phase, a heuristic algorithm is used to schedule a combined workflow. During the dynamic scheduling phase, the system seeks to provision the necessary resources and execute tasks based on the static plan. However, adjustments may be made to the initial schedule to account for uncertainties in the cloud environment. Similarly, Zhou et al. (Zhou et al., 2017) introduced a comparable framework with an offline optimization stage using breadth-first search or A* algorithms to create a static schedule, followed by an online adaptation stage that refines the primary schedule using runtime information about cloud resources and workflow executions. Figure 6(a) illustrates the distribution of articles across each category. It is evident that immediate dynamic scheduling is the most commonly used approach.
@@
Figure 6. Number of articles for different scheduling types and algorithms
4.2.7. Scheduling Algorithm
@@
The MWSP in a cloud environment is a well-known NP-hard problem (Ullman, 1975). Various algorithms have been employed by researchers to address this issue, with heuristic methods being the most commonly applied. The term heuristic refers to a method based on experience or judgment, designed to provide a solution that is close to optimal for a given problem (Silver, 2004). The majority of these studies have utilized the list heuristic to solve the MWSP (Chen et al., 2021; Zhou et al., 2018; Wang et al., 2021; Arabnejad et al., 2019; Xu et al., 2009; Ye et al., 2023). This heuristic involves two stages: task selection and resource mapping. In the task selection phase, each workflow task is assigned a rank according to a specific formula, after which the tasks are sorted and selected one by one based on their ranks. In the resource mapping phase, each selected task is assigned to a resource that optimizes the scheduling objective. Another commonly used heuristic in this field is the clustering heuristic (Li et al., 2024b; Ahmad et al., 2022), where tasks of the workflow are grouped into clusters, and each cluster is scheduled on a single resource to minimize data transmission time between tasks (in direct transmission mode). Other heuristic algorithms, such as A*, have also been explored by researchers (Zhou et al., 2016; Zhou et al., 2017).
@@
Meta-heuristic methods are another popular class of algorithms for task scheduling. Meta-heuristics are higher-level heuristics that are problem-independent and provide generic frameworks for finding near-optimal solutions to optimization problems (Abdel-Basset et al., 2018). Li et al. (Li et al., 2022c) proposed the Multi-swarm Co-evolution-based Hybrid Intelligent Optimization (MCHO) algorithm, which aims to minimize makespan and cost while meeting the deadline of each workflow. They utilized the Particle Swarm Optimization (PSO) algorithm, incorporating three different swarms: the first and second swarms focus on optimizing makespan and cost, respectively, while the third swarm is introduced to balance the two objectives. Additionally, they employed Genetic Algorithms (GA) as an elite enhancement strategy for each swarm, aiming to exploit elite individuals and increase the likelihood of finding more non-dominated solutions. Non-dominated Sorting Genetic Algorithm III (NSGA-III) is also applied to address the MWSP as a prominent multi-objective meta-heuristic method (Li et al., 2024b, c). Li et al. (Li et al., 2024b) first introduced a clustering heuristic called Depth-First-Search considering Critical Subsequent Task (DFS-CST) to partition each workflow into clusters and then applied NSGA-III to assign resources to each cluster, aiming to minimize three objectives simultaneously: execution time, execution cost, and unfairness.
@@
Zhou et al. (Zhou et al., 2024) introduced the Evolutionary Multitasking Constrained Optimization-Adaptive Double-level knowledge Transfer (EMCO-ADT) method for deadline-constrained multi-workflow scheduling. They treat each workflow scheduling problem as an individual ”task” that can be addressed using Evolutionary Algorithms (EAs). To enhance the performance of multi-workflow scheduling, they incorporate cross-task knowledge transfer between independent workflow schedules. This approach utilizes the Evolutionary Multi-Task Optimization (EMTO) paradigm, which facilitates knowledge extraction and transfer across distinct tasks. Qin and Shao (Qin and Shao, 2025) proposed a bi-population co-evolutionary algorithm for multi-objective, deadline-constrained multi-workflow scheduling. Their method involves two sub-populations: a deadline-aware sub-population that focuses on finding feasible schedules that meet deadlines, and a multi-objective sub-population that aims to minimize both makespan and cost. After each iteration, a bi-directional exchange of schedules occurs between the two sub-populations, promoting the diversity of each population.
@@
Other meta-heuristic algorithms, such as the Owl Search Algorithm (OSA) (Li et al., 2022d), Poor and Rich Optimization (PRO) algorithm (Li et al., 2024d), the Snake Optimizer (SO) (Li et al., 2023b), and the Gradient-Based Optimizer (GBO) (Li et al., 2024a), have also been applied to address the MWSP. Additionally, three studies have utilized Genetic Programming Hyper-Heuristic (GPHH) to generate heuristic rules for multi-workflow scheduling (Fan et al., 2023b; Sun et al., 2024; Escott et al., 2022). In this approach, each heuristic is represented as a syntax tree (GP tree), where the leaves are terminal nodes corresponding to problem parameters such as task execution time or a VM’s execution speed or cost, and the internal nodes represent function nodes, such as addition or maximum. Each heuristic is treated as an individual evolved by GPHH, and ultimately, the best-performing heuristic(s) is selected for scheduling.
@@
Recently, Reinforcement Learning (RL) has been applied to solve the MWSP. RL is a machine learning approach where an agent learns through interaction with an environment, making decisions based on feedback from its actions and experiences. The agent selects actions based on its policy and the environment’s state, receives rewards or penalties as feedback, and adjusts its policy accordingly. Li et al. (Li et al., 2022a) use a Double Deep Q-Network (DDQN) to minimize both makespan and cost in a multi-workflow scheduling context. They employ a two-level scheduling strategy: at the first level, a ready task is selected for scheduling, and at the second level, a suitable VM is chosen for that task. Each level features two sub-agents: a time sub-agent and a cost sub-agent, aimed at optimizing time and cost, respectively. After training, the model achieves an optimal balance between makespan and cost after 400 episodes. Pan and Wei (Pan and Wei, 2024) proposed a similar dynamic scheduling algorithm using a Deep Q-Network (DQN), but focusing solely on the VM selection phase. In another related study, Huang et al. (Huang et al., 2022) utilize a Deep Neural Network (DNN) for VM selection of the ready tasks, combined with an evolutionary algorithm to optimize the DNN’s parameters.
@@
Lin et al. (Lin et al., 2024) also employed RL for workflow task scheduling across a combination of on-demand and spot instances, though in a different way. For task scheduling, they used a DNN that takes an 11-dimensional vector of the ready task’s features as input. Instead of selecting a specific VM for task execution, the output is a meta-policy. The meta-policies are as follows: Util-priority: schedule the task on the available spot instance with the least available resources; Avl-priority: schedule the task on newer spot instances; To-spot: launch a new spot instance and schedule the task on it; To-on-demand: schedule the task on an idle available on-demand instance, or if none are available, launch a new on-demand instance; Wait: defer the task’s execution until the next scheduling period. While all the aforementioned works proposed dynamic scheduling algorithms, Zhang et al. (Zhang et al., 2023) introduced a static scheduling algorithm for multi-workflows. Their approach first utilizes a GA meta-heuristic to determine the optimal ordering of the workflow’s tasks, and then employs a DQN agent to assign tasks to VMs.

### v2

@@ 4.2.3 Task Concurrency
Task concurrency defines the number of tasks that can be executed concurrently on the same resource. Most studies adopt a single mode, where only one task is scheduled on each resource at a time [35, 161]. However, some studies explore the multiple mode, which permits the concurrent execution of several tasks on the same resource. For instance, in [159, 139], brokers rent large VMs with a high number of CPU cores and significant memory, bundling tasks together and scheduling them on these VMs as containers. Similarly, Silva et al. [153] utilize multi-threading on a single node to enable the simultaneous execution of multiple tasks.
4.2.4 Data Transmission
@@
Data transmission outlines how output data from parent (predecessor) tasks is delivered to child (successor) tasks. In the direct (or peer-to-peer) approach, data is transferred straight from the resource executing the predecessor task to the resource assigned to the successor task [33, 17]. This method requires synchronous communication, meaning the successor task must already be scheduled, and its resource allocation must be determined beforehand. A notable advantage of this approach is that data transfer time between tasks scheduled on the same resource becomes negligible. However, it may not be ideal for scheduling strategies that delay the execution of successor tasks. In contrast, the indirect (or shared) method involves the predecessor task storing its output data in shared storage [130, 129]. The successor task can then retrieve this data asynchronously, enabling more flexibility in task scheduling.
4.2.5 Scheduling Criteria
The most frequently addressed criteria are the execution time of workflows and the rental cost of resources, which appear in nearly all studies either as explicit optimization objectives or as constraints (such as deadlines or budgets). In addition, several studies highlight the importance of maximizing rented resource utilization [111, 36, 81], considering it a key measure of the efficiency and effectiveness with which scheduling algorithms leverage available cloud resources.
Another frequently studied objective is the minimization of energy consumption. Although energy efficiency is typically considered more of a provider-side concern, a number of studies have explored mechanisms that can be implemented at the broker side to reduce energy usage. One line of research proposes dynamically adjusting the CPU frequencies of leased resources through dynamic voltage and frequency scaling (DVFS) technology [161, 160, 104], thereby lowering energy consumption during execution. Another approach [165] employs consolidation, where tasks are assigned to the minimum number of resources to maximize utilization. In this strategy, tasks are preferentially scheduled on highly utilized resources, enabling underutilized hosts to become idle and transition into sleep (hibernate) mode, ultimately saving energy. Along similar lines, [30] models the energy consumption of each resource as a linear function of its utilization and seeks to maintain utilization at optimized levels to minimize total energy usage. In hybrid cloud environments, where both private and public clouds are used, researchers often emphasize reducing energy consumption in the private cloud while disregarding the rental costs incurred in the public cloud [104, 156].
@@ 4.2.5 Scheduling Criteria
Fairness has also emerged as an important criterion, though its definition varies across the literature. Some studies [35, 184] treat fairness as a resource-centric property aimed at ensuring balanced workload distribution among resources. In contrast, other works [97, 98, 131, 147] adopt a user-centric perspective, emphasizing equitable resource allocation among workflows. For instance, [97] introduces the concept of a workflow’s loss rate, which combines two factors: slowdown (the ratio of a workflow’s completion time when scheduled alongside others to its completion time in isolation) and overspending (the cost ratio under the same conditions). Fairness is then defined as minimizing disparities in loss rates across workflows. Similarly, Yang et al. [147] propose slowdown variation as a fairness metric, measuring the sum of absolute differences between each workflow’s slowdown and the average slowdown across all workflows.
Three studies explicitly addressed privacy in hybrid private/public clouds [156, 148, 67]. They define privacy-sensitive tasks as those containing private information, such as financial or medical data, which must be executed on the local private cloud. Ordinary tasks, by contrast, may be executed on either public or private resources. Reference [148] further introduced an intermediate category of tasks that can run on private resources or on a restricted subset of public resources.
Security has also been examined in several studies [8, 138, 7, 67]. These works converge on a similar definition, encompassing three core services: authentication, integrity (via hash algorithms), and confidentiality (via cryptographic algorithms). Together, these services aim to defend against major security threats such as identity spoofing, data alteration, and unauthorized access or snooping. In this framework, each workflow task is assigned a security demand (or level) for each service, ranging from 0 to 1, with higher values reflecting greater sensitivity. Scheduling decisions are then made such that each task is executed on a resource or service whose security level meets or exceeds its requirements.
Another study investigated trust [105] as a QoS requirement from the perspectives of both users and CSPs. The authors define trust as the subjective judgment of the trustor, based on personal experience and relevant knowledge, including authenticity, integrity, reliability, and stability. For a user to transact with a CSP, the decision process begins by checking their history of direct interactions with that CSP. If sufficient past transactions exist, the user relies solely on their computed direct trust. Otherwise, the user broadcasts a request for recommendations to other trusted entities and calculates an integrated trust value by combining limited direct trust with recommended trust obtained from others’ opinions. This integrated value forms the basis for the final decision. CSPs, in contrast, adopt a simpler mechanism: they maintain a blocklist of known malicious entities and apply the straightforward rule that “everyone is allowed to request service unless they are on the blocklist.”
Reliability has been considered in two studies, where it is defined as the probability of completing a task without failure on a given resource. In [161], reliability is computed from the resource’s fault rate and operating frequency, and the authors employ task duplication to improve the reliability of tasks deemed less reliable. In contrast, [141] explores the joint provisioning of spot and on-demand instances. In this model, the reliability of executing a task on a spot instance is estimated through a cumulative distribution function, which approaches unity as the task’s finish time moves further from the instance’s launch time—reflecting the observation that freshly launched spot instances have lower failure probabilities than older ones. The authors then aim to strike a balance among time, cost, and reliability when scheduling workflow tasks.
4.2.6 Scheduling Type
@@
Workflow scheduling can be classified into two primary types: static and dynamic [136]. Static scheduling involves assigning tasks to resources prior to the start of workflow execution. This approach typically produces high-quality solutions since it considers crucial structural aspects of the workflow, such as the critical path. However, static scheduling is limited by its reliance on precise runtime estimations and its inability to adapt to unforeseen changes in resource availability, making it less suitable for the uncertain nature of cloud environments. On the other hand, dynamic scheduling makes decisions during runtime, assigning tasks as they become ready—that is, once all their predecessor tasks have been completed. This enables dynamic methods to respond effectively to changes and uncertainties in the cloud environment. Nevertheless, these algorithms often make greedy, task-specific decisions, lacking a global perspective on the entire workflow. Consequently, the solutions generated by dynamic approaches are generally less optimal compared to those of static methods. To address the shortcomings of both approaches, hybrid strategies have been proposed. These methods begin by creating a static execution plan for the workflow and then apply dynamic adjustments during runtime to account for environmental uncertainties. This combination leverages the strengths of both static and dynamic scheduling to produce more robust solutions.
@@
In static multiple workflow scheduling, two primary approaches have been used: one-by-one and combined. In the one-by-one method, workflows are scheduled sequentially, based on priority, while considering resource availability [131, 132]. For example, Rezaeian et al. [131] apply the one-by-one approach using four different workflow selection methods: Ordered, which schedules workflows based on a predefined priority; round-robin, which iterates over workflows, scheduling one task from each workflow at a time; weighted round-robin, where workflows are selected according to a weighted formula before scheduling a task; and unfairness heuristic, which selects workflows based on their level of unfairness. In contrast, the combined approach merges all workflows into a single large workflow before scheduling it on available resources. References [8, 5] use a similar approach, where they break down each workflow into levels based on task depth. Then, they create partitions where each partition contains tasks at the same level across different workflows. Other studies merge workflows by adding a dummy entry task to connect the input workflows’ actual entry tasks, and a dummy exit task to connect the workflows’ exit tasks [163, 128, 195].
@@
Two strategies are commonly used for dynamic multiple workflow scheduling: immediate and periodic. In the immediate strategy, tasks are scheduled as soon as they become ready. However, this approach faces two challenges: first, the runtime overhead caused by executing the scheduling algorithm for each task as it becomes ready, and second, the greedy decision-making process for each task without considering the future tasks and their potential impact on the current scheduling. To address these issues, the periodic strategy was introduced. This method divides time into intervals, collecting all ready tasks within each interval and scheduling them all at once at the end of the interval. While this approach resolves the problems associated with immediate scheduling, it also introduces a delay, causing ready tasks to wait until the end of the interval, potentially missing deadlines. Most previous work in dynamic scheduling adopts the immediate strategy [110, 112, 130], although some utilize the periodic strategy [159, 109]. Chen et al. [34] introduced a hybrid strategy, where they use periodic scheduling but select tasks at the start of each interval that are predicted to have their earliest start time within the current interval. These tasks are scheduled on available resources ahead of time.
@@
Four studies have employed a combination of both static and dynamic scheduling types [163, 196, 172, 166]. Taheri et al. [163] proposed a framework consisting of two phases: static planning and dynamic scheduling. In the static planning phase, a heuristic algorithm is used to schedule a combined workflow. During the dynamic scheduling phase, the system seeks to provision the necessary resources and execute tasks based on the static plan. However, adjustments may be made to the initial schedule to account for uncertainties in the cloud environment. Similarly, Zhou et al. [196] introduced a comparable framework with an offline optimization stage using breadth-first search or A* algorithms to create a static schedule, followed by an online adaptation stage that refines the primary schedule using runtime information about cloud resources and workflow executions. Finally, Toporkov et al. [166] perform a static analysis of each incoming workflow using the Critical Jobs’ Method (CJM) to determine feasible execution windows for all tasks. These windows are then used to group tasks into independent execution batches that respect data dependencies. At runtime, as each batch approaches its planned start time, a dynamic assignment process selects and provisions suitable VMs by solving a matching problem.
Figure 6(a) illustrates the distribution of articles across each category. It is evident that immediate dynamic scheduling is the most commonly used approach.
@@
Figure 6: Number of articles for different scheduling types and algorithms
4.2.7 Scheduling Algorithm
@@
The MWSP in a cloud environment is a well-known NP-hard problem [167]. Various algorithms have been employed by researchers to address this issue, with heuristic methods being the most commonly applied. The term heuristic refers to a method based on experience or judgment, designed to provide a solution that is close to optimal for a given problem [154]. The majority of these studies have utilized the list heuristic to solve the MWSP [35, 200, 171, 17, 188, 190, 20]. This heuristic involves two stages: task selection and resource mapping. In the task selection phase, each workflow task is assigned a rank according to a specific formula, after which the tasks are sorted and selected one by one based on their ranks. In the resource mapping phase, each selected task is assigned to a resource that optimizes the scheduling objective. Another commonly used heuristic in this field is the clustering heuristic [97, 5], where tasks of the workflow are grouped into clusters, and each cluster is scheduled on a single resource to minimize data transmission time between tasks (in direct transmission mode). Other heuristic algorithms, such as A*, have also been explored by researchers [197, 196].
@@
Meta-heuristic methods are another popular class of algorithms for task scheduling. Meta-heuristics are higher-level heuristics that are problem-independent and provide generic frameworks for finding near-optimal solutions to optimization problems [1]. Li et al. [102] proposed the Multi-swarm Co-evolution-based Hybrid Intelligent Optimization (MCHO) algorithm, which aims to minimize makespan and cost while meeting the deadline of each workflow. They utilized the Particle Swarm Optimization (PSO) algorithm, incorporating three different swarms: the first and second swarms focus on optimizing makespan and cost, respectively, while the third swarm is introduced to balance the two objectives. Additionally, they employed Genetic Algorithms (GA) as an elite enhancement strategy for each swarm, aiming to exploit elite individuals and increase the likelihood of finding more non-dominated solutions. Non-dominated Sorting Genetic Algorithm III (NSGA-III) is also applied to address the MWSP as a prominent multi-objective meta-heuristic method [97, 98]. Li et al. [97] first introduced a clustering heuristic called Depth-First-Search considering Critical Subsequent Task (DFS-CST) to partition each workflow into clusters and then applied NSGA-III to assign resources to each cluster, aiming to minimize three objectives simultaneously: execution time, execution cost, and unfairness.
@@
Qin and Shao [128] proposed a bi-population co-evolutionary algorithm for multi-objective, deadline-constrained multi-workflow scheduling. Their method involves two sub-populations: a deadline-aware sub-population that focuses on finding feasible schedules that meet deadlines, and a multi-objective sub-population that aims to minimize both makespan and cost. After each iteration, a bi-directional exchange of schedules occurs between the two sub-populations, promoting the diversity of each population. Other meta-heuristic algorithms, such as the Simulated Annealing (SA) [187], Owl Search Algorithm (OSA) [104], Poor and Rich Optimization (PRO) algorithm [101], the Snake Optimizer (SO) [103], the Gradient-Based Optimizer (GBO) [99], and the Linear Congruential Lyrebird Optimization Algorithm (LC-LOA) [88] have also been applied to address the MWSP.
@@
Hyper-heuristics have also been applied to address MWSP. Unlike meta-heuristics, which directly search the solution space, hyper-heuristics operate at a higher level by selecting, combining, generating, and sequencing problem-specific heuristics from a predefined set of heuristics and/or meta-heuristics [48]. Their search is conducted in the space of heuristics rather than the space of problem solutions. In particular, three studies have employed Genetic Programming Hyper-Heuristics (GPHH) to evolve heuristic rules for multi-workflow scheduling [54, 157, 52]. In this approach, each heuristic is represented as a syntax tree (also known as a GP tree), where the terminal nodes correspond to problem parameters—such as task execution time, or the execution speed and cost of a VM—while the internal nodes represent function operators, such as addition or maximum. Each tree represents a candidate heuristic evolved by GPHH, and the most effective heuristic(s) are ultimately selected for scheduling.
@@
Recently, Reinforcement Learning (RL) has been applied to solve the MWSP. RL is a machine learning approach where an agent learns through interaction with an environment, making decisions based on feedback from its actions and experiences. The agent selects actions based on its policy and the environment’s state, receives rewards or penalties as feedback, and adjusts its policy accordingly. Li et al. [100] use a Double Deep Q-Network (DDQN) to minimize both makespan and cost in a multi-workflow scheduling context. They employ a two-level scheduling strategy: at the first level, a ready task is selected for scheduling, and at the second level, a suitable VM is chosen for that task. Each level features two sub-agents: a time sub-agent and a cost sub-agent, aimed at optimizing time and cost, respectively. After training, the model achieves an optimal balance between makespan and cost after 400 episodes. Pan and Wei [125] proposed a similar dynamic scheduling algorithm using a Deep Q-Network (DQN), but focusing solely on the VM selection phase. In another related study, Huang et al. [69] utilize a Deep Neural Network (DNN) for VM selection of the ready tasks, combined with an evolutionary algorithm to optimize the DNN’s parameters.
Chen et al. [31] proposed Arbitrary-Freedom Adaptive Scheduling (AFAS), which also employs Deep Reinforcement Learning (DRL) to prioritize ready tasks and generate a scheduling sequence for mapping them to available resources. In their approach, workflow tasks are categorized into two types: routine tasks, which appear frequently with predictable patterns and resource requirements, and nonroutine tasks, which occur irregularly and exhibit variable resource demands. The proposed adaptive scheduling algorithm combines a heuristic-based scheduling method for handling nonroutine tasks with a DRL-based intelligent scheduling strategy tailored for routine tasks. In another study, He et al. [67] first employed a Q-learning approach to assign priorities to the ready tasks, and then used a DRL method to map the prioritized tasks to available VMs.
@@
Lin et al. [109] also employed RL for workflow task scheduling across a combination of on-demand and spot instances, though in a different way. For task scheduling, they used a DNN that takes an 11-dimensional vector of the ready task’s features as input. Instead of selecting a specific VM for task execution, the output is a meta-policy. The meta-policies are as follows: Util-priority: schedule the task on the available spot instance with the least available resources; Avl-priority: schedule the task on newer spot instances; To-spot: launch a new spot instance and schedule the task on it; To-on-demand: schedule the task on an idle available on-demand instance, or if none are available, launch a new on-demand instance; Wait: defer the task’s execution until the next scheduling period. While all the RL-based approaches focus on dynamic scheduling, Zhang et al. [191] proposed a static scheduling algorithm for handling multiple workflows. Their approach first utilizes GA to determine the optimal ordering of the workflow’s tasks, and then employs a DQN agent to assign the whole workflow to a VM instance.
To prepare richer feature vectors for workflow tasks and resources, several studies have employed Graph Neural Networks (GNNs) [199]. GNNs are a class of deep learning models designed to operate on graph-structured data (consisting of nodes and edges). They work by aggregating information from a node’s neighbors and updating the node’s representation based on its own features and the features of its neighboring nodes. This mechanism allows GNNs to capture both the attributes of individual nodes and the overall graph topology. The output of a GNN is typically an embedding—a numerical vector representation of a node (or edge or entire graph)—that encapsulates its properties and structural context. These embeddings can then be used as input features to a reinforcement learning-based method.
For example, Wang et al. [177] employed a Graph Attention Network (GAT)—a specific type of GNN—to encode the features of each ready task along with its neighboring tasks. They applied a similar strategy to generate embeddings for each VM in the network. The concatenation of these task and VM embeddings formed the input to a DRL algorithm, which was then used to schedule ready tasks on the available resources. Along the same line, Huang et al. [70] also integrated GAT with DRL, but their objective was specifically to minimize execution costs while meeting deadline. Finally, Chandrasiri and Meedeniya [30] proposed a method that constructs a heterogeneous graph whose nodes include both workflow tasks and VMs. A GNN is then used to compute embeddings for this combined graph. These embeddings serve as inputs to a DRL-based scheduler that makes task-to-VM assignment decisions with the dual objective of minimizing both makespan and energy consumption.
Meta-learning, often described as learning to learn, is a paradigm where algorithms are trained not just to solve individual learning problems but also to acquire the ability to adapt quickly to new ones [170]. Instead of starting from scratch for each new scheduling scenario, a meta-learner captures transferable knowledge across related scheduling problems and uses it to accelerate adaptation to unseen workflows. In the context of multiple workflow scheduling, this means the scheduler can exploit similarities between workflows—such as recurring DAG structures, resource contention patterns, or typical deadline/cost trade-offs—to refine strategies more efficiently.
One branch of meta-learning is Evolutionary Multi-Task Optimization (EMTO), which evolves scheduling heuristics or policies simultaneously across several scheduling problems. In this way, insights gained in optimizing one workflow (e.g., dealing with long critical paths or VM bottlenecks) can be transferred across workflows, reducing search costs and improving overall efficiency. This knowledge transfer between scheduling problems is particularly valuable in WaaS environments, where workflows often arrive with diverse characteristics but share common structural or performance patterns. Zhou et al. [198] employed EMTO to address deadline-constrained multi-workflow scheduling, while Zhang et al. [194] applied a similar approach to develop a scheduling algorithm that simultaneously minimizes makespan, cost, and energy consumption. Another widely studied meta-learning approach is meta-reinforcement learning (meta-RL) [23], which has been applied to the scheduling of multiple workflows in hybrid edge–cloud environments [39, 41]. However, its application in cloud-only settings has not yet been investigated, making it a promising direction for future research (see Section 5.1).